OpenStudy (anonymous):

Peanuts worth $4/kg were mixed with almonds worth $6/kg to produce 12 kg of nuts worth $62. Represent the unknowns above using the variables x and y.

OpenStudy (anonymous):

4x+6y=12*62 x+y=12

OpenStudy (anonymous):

x and y represent amount of peanuts in Kg

OpenStudy (anonymous):

Write an equation that models the relationships between the weights of the nuts in the mixture.? is it the same as you said ? 4x+6y=12*62 x+y=12

OpenStudy (anonymous):

look same to me